OpenStudy (anonymous):

what is the asymptotes for y=3x-1/√9x^2-1

OpenStudy (anonymous):

You will have a vertical asymptote at the place where the denominator is equal to zero, hence the equation is not defined.

OpenStudy (anonymous):

for the remaining asymptotes, you can see how this equations acts as the limit of x approaches infinity.

OpenStudy (anonymous):

but the answer is y=+1 and -1,x=+1/3 and -1/3

OpenStudy (anonymous):

\[ \sqrt{9x^2-1}=0 \\ 9x^2=1 \\x= \pm \frac{1}{3} \]

OpenStudy (anonymous):

for horizontal asymptote?

OpenStudy (anonymous):

Sorry I had to disappear for a moment.

OpenStudy (anonymous):

You know how horizontal asymptotes can be found? \[ \lim_{x \to \infty} \frac{3x-1}{\sqrt{9x^2-1}} \]

OpenStudy (anonymous):

the annoying part is that you only have a radical in the denominator, but you can easily help yourself to get rid of this expression. \[\pm (3x-1) = (\sqrt{3x-1})^2 \] See where this is going?

OpenStudy (anonymous):

\[ \sqrt{\frac{9x^2-6x+1}{9x^2-1}} \] Equal grade between numerator and denominator, hence the horizontal limit is: \[ y= \pm 1 \]